# BELLA: Black-box model Explanations by Local Linear Approximations

### v2

@@ 5 Experiments
Datasets. We performed experiments on datasets from two standard repositories [Dua and Graff, 2017, Romano et al., 2021] (shown in Table 1). Among them is also a high-dimensional dataset, Superconductivity, with 81 features. All categorical features have been one-hot encoded and all numerical features have been standardized. We draw a random 10% of each dataset as testing data. To show that BELLA works with different families of models, we trained a random forest (with 100 trees), and a neural network (with one hidden layer with 500 nodes) as black-box models. Since the results do not differ much, we show only experiments with the neural network here, while the experiments with the random forest are in the supplementary material.
BELLA. Our method is implemented in Python. We set the step size to 10%. For the black-box models, we use the implementations of scikit-learn [Pedregosa et al., 2011]. All experiments are run on a Fedora Linux (release 38) computer with an Intel(R) Xeon(R) v4 @ 2.20GHz CPU, a memory of 64 GB, and Python 3.9. All code and the data for BELLA and the experiments is available on Github at https://github.com/nedRad88/BELLA.
@@ 5.1 Experimental results
Factual Counterf.
Dataset LIME MAPLE BELLA SHAP BELLA
Auto MPG 2.99 $\pm$0.830 0.86 $\pm$0.270 1.45 $\pm$0.390 0.00 5.69 $\pm$1.210
Bike 579.76 $\pm$24.73 75.42 $\pm$6.710 224.70 $\pm$12.90 0.00 1053.64 $\pm$4.25
Concrete 10.61 $\pm$1.410 2.13 $\pm$0.290 4.87 $\pm$0.670 0.00 10.73 $\pm$1.86
Servo 0.75 $\pm$0.320 0.21 $\pm$0.100 0.59 $\pm$0.260 0.00 1.12 $\pm$0.260
Electrical 0.02 $\pm$0.002 0.01 $\pm$0.001 0.02 $\pm$0.002 0.00 0.02 $\pm$0.003
Supercond. 23.17 $\pm$0.697 1.05 $\pm$0.029 14.24 $\pm$0.434 0.00 57.50 $\pm$1.150
White Wine 0.36 $\pm$0.030 0.17 $\pm$0.010 0.29 $\pm$0.020 0.00 1.53 $\pm$0.090
Real Estate 4.97 $\pm$1.170 1.48 $\pm$0.540 2.01 $\pm$0.730 0.00 1.76 $\pm$1.700
Wind 2.52 $\pm$0.382 1.15 $\pm$0.173 1.69 $\pm$0.247 0.00 5.43 $\pm$0.753
CPU Activity 16.30 $\pm$1.620 0.81 $\pm$0.060 1.18 $\pm$0.110 0.00 2.24 $\pm$0.220
Echocard. 3.02 $\pm$0.046 1.82 $\pm$0.031 2.84 $\pm$0.049 0.00 6.86 $\pm$0.101
Iranian Churn 172.13 $\pm$23.52 4.04 $\pm$0.970 24.40 $\pm$7.950 0.00 30.49 $\pm$3.770
Norm. avg. 0.10 $\pm$0.014 0.02 $\pm$0.004 0.05 $\pm$0.008 0.00 0.14 $\pm$0.015
@@
Table 3: Generality comparison (- larger is better)
@@
Generality is measured by the number of data points to which the explanation applies (as a percentage of all data points in the training set). One could give a hypercube for the size of the neighborhood, but it is arguably the number of data points (and not the size of a potentially sparse hypercube) that conveys the significance of the explanation. Thus, for BELLA, we simply return the size of the neighborhood. For MAPLE we return the number of data points that have weights larger than 00. For LIME, an explanation comes with the range of values for each feature. We count the number of data points that fall into this range. The results are shown in Table 3. For SHAP, the size of the neighborhood is always 0. This is because SHAP provides feature contributions that are specific for the given data point, and there is no way to apply these explanations to other data points. LIME’s explanations are more general, and MAPLE’s explanations even more. Still, they are vastly less general than the explanations of BELLA.
@@
$\textit{robustness}=1-\frac{1}{n}\sum_{i=1}^{n}\frac{|\beta_{1i}-\beta_{2i}|}{|\beta_{1i}|+|\beta_{2i}|}.$ (5)
Here, $n$ is the number of features, and $\beta_{1i}$ and $\beta_{2i}$ are the weights of feature $i$ in the first and second explanation, respectively. Robustness is in the range of $[0,1]$, with 1 indicating that two explanations are identical. We compute explanations for each data point in the test set, and compute robustness wrt. a smaller set of 5 closest neighbors, and a larger set of 20 closest neighbors (Table 5). As expected, all methods are a bit less robust when the set of neighbors is larger, but otherwise the results are very similar: LIME samples $5000$ data points to create a synthetic neighborhood. Thus, LIME can perform better than our approach on datasets that have fewer observations. However, in the vast majority of cases, as well as on average, BELLA outperforms LIME. BELLA also outperforms SHAP by a wide margin. This is because SHAP’s explanations are tailored for a single data point. BELLA also outperforms MAPLE. This is because the crisp neighborhood of BELLA provides much more robust explanations than MAPLE’s weighted neighborhood.
@@
From Tables 234, and 5, we can see that at the same level of simplicity, BELLA provides more general, more robust, and more accurate explanations than LIME. BELLA provides less accurate explanations than SHAP and MAPLE, but at the same time, BELLA’s explanations are more general, more robust, and vastly simpler. The results when the black-box model is a random forest are shown in the appendix, and they do not differ much.
@@
Runtime of all methods is shown in Table 6. On average, MAPLE and SHAP are extraordinarily fast, and LIME is slower. BELLA is, on average, 5$\times$ slower than LIME. This is due mainly to a single dataset, Superconductivity, which has a very large number of features, and on which BELLA is 11$\times$ slower than LIME. This is because, different from our competitors, BELLA is deterministic. Our method thus has to explore the whole local space. At the same time, BELLA runs in the same order of magnitude of time as LIME on average, and it remains thus competitive. On the median, BELLA is even faster than LIME.
Counterfactuality is a desideratum that only BELLA, and neither LIME nor SHAP nor MAPLE fulfills. To evaluate the quality of BELLA’s counterfactual explanations, we measure the fidelity of an explanation wrt. a reference value. For each data point $x$ with its target value $y$, we set the reference values to $y_{\textit{ref}}=y\pm 0.3\times|y_{\textit{max}}-y_{\textit{min}}|$, so as to be different, but not unrealistically far away from $y$. The last column of Table 2 shows the RMSE of the counterfactual explanations (as well as of the factual explanations by BELLA and LIME for reference). We see that the counterfactual explanations of BELLA are often of similar fidelity as its factual explanations. Even in the cases where the error of counterfactual explanations is an order of magnitude larger, it is still lower or comparable to the error of the factual-only explanations provided by LIME.
Other desiderata outlined in Section 3 were determinism, and verifiability (the possibility to compute the explained value from the feature values). SHAP offers none of these. Neither does LIME. While both SHAP and LIME compute linear models with feature weights, these models are not verifiable in our sense: There is no way that the user can insert the feature values of a neighboring point into these models and obtain an explained value. This is because the linear models do not operate in the original input feature space. Only MAPLE offers this verifiability. However, it relies on randomization and provides no counterfactuality. BELLA is thus the only approach that delivers deterministic, verifiable, and both factual and counterfactual explanations.
@@ 6 Limitations
BELLA is a domain-independent method to produce deterministic post-hoc local explanations of tabular regression datasets. This scope entails some inherent limitations: First, as a generalist method, BELLA does not take into account domain-specific peculiarities, such as domain-specific weighting, dynamic scaling of feature distances, or domain-specific distance functions. BELLA also currently does not take into account domain-specific lists of features that can or cannot be modified, so that it may produce de facto impossible counterfactuals (as in: “To avoid a cancer diagnostic, modify your age to be 10 years younger”). It may thus prove necessary to adapt BELLA to specific domains. While such adaptations are certainly possible, they are out of the scope of the present paper. Second, BELLA aims at simple and verifiable (and ultimately linear) explanations. With this, BELLA may fail to capture intricate, higher-order interactions of features. This problem is part of the intrinsic trade-off between simplicity and accuracy, which any post-hoc explainability approach finds itself in. Third, BELLA has to explore the whole neighborhood space without sampling, because it is deterministic. This incurs a high algorithmic cost. In future work, this cost could be alleviated by using more efficient data structures (such as KD-trees). One could even give up on the desideratum of determinism, and resort to sampling. Such advances are, likewise, left for future work.
There are also a number of ethical considerations: First, BELLA does not correct for biases in the data or the model. If the data contains biased samples, discriminating features, or an otherwise unprofessional selection of features or data points, these characteristics will be mirrored in BELLA’s explanations. We believe, however, that this is a feature rather than a bug: BELLA’s explanations make such biases visible to the user. It would be disastrous to correct them and thus convey the impression that the model does not have them. A second concern is the problem of overtrust, where users mistakenly take BELLA’s local explanation for a global correlation. That is an issue inherent to all local approaches, which would have to be countered by educating the user before using the approach. Finally, there is always the possibility that users “game” the model without genuinely improving underlying attributes (e.g., artificially adjusting specific features to increase a credit score). BELLA’s counterfactual explanations might then recommend such tweaks also to other users.
@@ 7 Conclusion
We have presented BELLA, an approach to provide post-hoc local explanations for any regression black-box model, or indeed any static tabular dataset with a numeric variable to be explained. BELLA is deterministic, and can provide both factual and counterfactual explanations. BELLA’s objective function ensures accurate, general, robust, and simple explanations. Detailed experiments show that BELLA outperforms state-of-the-art approaches on these desiderata, often by a wide margin.
Future work could investigate methods to improve the speed of BELLA. One possibility is to use other data structures to speed up the neighborhood search. Another possibility is to give up determinism and resort to sampling to further speed up the algorithm. Finally, one could investigate even if BELLA could replace black box models for making a prediction in the first place, following positive experiences with linear models elsewhere [Ismail et al., 2022].
We hope that our work can open the door to research along this line and others, and ultimately make machine learning models more interpretable.
@@ Appendix A Experiments with Random Forest as black-box model
Factual Counterf.
Dataset LIME MAPLE BELLA SHAP BELLA
Auto MPG 1.57 $\pm$0.470 0.95 $\pm$0.272 1.52 $\pm$0.412 0.00 2.97 $\pm$0.880
Bike 338.50 $\pm$20.92 61.23 $\pm$5.560 228.14 $\pm$14.48 0.00 1057.2 $\pm$3.782
Concrete 5.78 $\pm$0.844 2.22 $\pm$0.341 5.11 $\pm$0.681 0.00 4.25 $\pm$0.942
Servo 0.45 $\pm$0.251 0.18 $\pm$0.152 0.45 $\pm$0.211 0.00 4.47 $\pm$1.17
Electrical 0.01 $\pm$0.002 0.00 $\pm$0.001 0.01 $\pm$0.002 0.00 0.00 $\pm$0.004
Supercond. 30.68 $\pm$1.407 0.96 $\pm$0.086 15.52 $\pm$0.481 0.00 57.50 $\pm$1.495
White Wine 0.30 $\pm$0.022 0.17 $\pm$0.014 0.28 $\pm$0.023 0.00 0.87 $\pm$0.063
Real Estate 5.19 $\pm$1.292 2.55 $\pm$0.962 5.04 $\pm$1.503 0.00 10.71 $\pm$2.632
Wind 1.40 $\pm$0.197 0.70 $\pm$0.128 1.17 $\pm$0.184 0.00 3.84 $\pm$0.492
CPU Activity 12.57 $\pm$1.110 0.63 $\pm$0.070 1.26 $\pm$0.010 0.00 2.54 $\pm$0.011
Echocard. 3.31 $\pm$0.510 1.64 $\pm$0.246 3.21 $\pm$0.532 0.00 16.87 $\pm$1.869
Iranian Churn 141.57 $\pm$20.44 9.46 $\pm$1.783 17.39 $\pm$2.923 0.00 320.59 $\pm$23.83
Norm. avg. 0.07 $\pm$0.010 0.02 $\pm$0.004 0.05 $\pm$0.008 0.00 0.18 $\pm$0.026
Table 8: Generality comparison (- larger is better)

### v3

@@ 5 Experiments
Datasets. We performed experiments on datasets from two standard repositories [Dua and Graff, 2017, Romano et al., 2021] (shown in Table 1). Among them is also a high-dimensional dataset, Superconductivity, with 81 features. All categorical features have been one-hot encoded and all numerical features have been standardized. We draw a random 10% of each dataset as testing data. To show that BELLA works with different families of models, we trained a random forest (with 100 trees), and a neural network (with one hidden layer with 500 nodes) as black-box models. Since the results do not differ much, we show only experiments with the neural network here, while the experiments with the random forest are in Appendix A.
BELLA. Our method is implemented in Python. We set the step size to 10%. For the black-box models, we use the implementations of scikit-learn [Pedregosa et al., 2011]. All experiments are run on a Fedora Linux (release 38) computer with an Intel(R) Xeon(R) v4 @ 2.20GHz CPU, a memory of 64 GB, and Python 3.9. All code and the data for BELLA and the experiments is available on Github (URL masked for anonymity).
@@ 5.1 Experimental results
Dataset LIME MAPLE BELLA SHAP
Auto MPG 2.99 $\pm$0.830 0.86 $\pm$0.270 1.45 $\pm$0.390 0.00
Bike 579.76 $\pm$24.73 75.42 $\pm$6.710 224.70 $\pm$12.90 0.00
Concrete 10.61 $\pm$1.410 2.13 $\pm$0.290 4.87 $\pm$0.670 0.00
Servo 0.75 $\pm$0.320 0.21 $\pm$0.100 0.59 $\pm$0.260 0.00
Electrical 0.02 $\pm$0.002 0.01 $\pm$0.001 0.02 $\pm$0.002 0.00
Supercond. 23.17 $\pm$0.697 1.05 $\pm$0.029 14.24 $\pm$0.434 0.00
White Wine 0.36 $\pm$0.030 0.17 $\pm$0.010 0.29 $\pm$0.020 0.00
Real Estate 4.97 $\pm$1.170 1.48 $\pm$0.540 2.01 $\pm$0.730 0.00
Wind 2.52 $\pm$0.382 1.15 $\pm$0.173 1.69 $\pm$0.247 0.00
CPU Activity 16.30 $\pm$1.620 0.81 $\pm$0.060 1.18 $\pm$0.110 0.00
Echocard. 3.02 $\pm$0.046 1.82 $\pm$0.031 2.84 $\pm$0.049 0.00
Iranian Churn 172.13 $\pm$23.52 4.04 $\pm$0.970 24.40 $\pm$7.950 0.00
Norm. avg. 0.10 $\pm$0.014 0.02 $\pm$0.004 0.05 $\pm$0.008 0.00
@@
Table 3: Generality comparison (% - larger is better)
@@
Generality is measured by the number of data points to which the explanation applies (as a percentage of all data points in the training set). One could give a hypercube for the size of the neighborhood, but it is arguably the number of data points (and not the size of a potentially sparse hypercube) that conveys the significance of the explanation. Thus, for BELLA, we simply return the size of the neighborhood. For MAPLE we return the number of data points that have weights larger than 0. For LIME, an explanation comes with the range of values for each feature. We count the number of data points that fall into this range. The results are shown in Table 3. For SHAP, the size of the neighborhood is always 0. This is because SHAP provides feature contributions that are specific for the given data point, and there is no way to apply these explanations to other data points. LIME’s explanations are more general, and MAPLE’s explanations even more. Still, they are vastly less general than the explanations of BELLA.
@@
$\textit{robustness}=1-\frac{1}{n}\sum_{i=1}^{n}\frac{|w_{1i}-w_{2i}|}{|w_{1i}|+|w_{2i}|}.$ (4)
Here, $n$ is the number of features, and $w_{1i}$ and $w_{2i}$ are the weights of feature $i$ in the first and second explanation, respectively. Robustness is in the range of $[0,1]$, with 1 indicating that two explanations are identical. We compute explanations for each data point in the test set, and compute robustness wrt. a smaller set of 5 closest neighbors, and a larger set of 20 closest neighbors (Table 5). As expected, all methods are a bit less robust when the set of neighbors is larger, but otherwise the results are very similar: LIME samples $5000$ data points to create a synthetic neighborhood. Thus, LIME can perform better than our approach on datasets that have fewer observations. However, in the vast majority of cases, as well as on average, BELLA outperforms LIME. BELLA also outperforms SHAP by a wide margin. This is because SHAP’s explanations are tailored for a single data point. BELLA also outperforms MAPLE. This is because the crisp neighborhood of BELLA provides much more robust explanations than MAPLE’s weighted neighborhood.
@@
From Tables 234, and 5, we can see that at the same level of simplicity, BELLA provides more general, more robust, and more accurate explanations than LIME. BELLA provides less accurate explanations than SHAP and MAPLE, but at the same time, BELLA’s explanations are more general, more robust, and vastly simpler. The results when the black-box model is a random forest are shown in Appendix A, and they do not differ much.
[…]
The other desiderata outlined in Section 3 were determinism and verifiability (the possibility to compute the explained value from the feature values). SHAP offers none of these. Neither does LIME. While both SHAP and LIME compute linear models with feature weights, these models are not verifiable in our sense: There is no way that the user can insert the feature values of a neighboring point into these models and obtain an explained value. This is because the linear models do not operate in the original input feature space. Only MAPLE offers this verifiability. However, it relies on randomization. BELLA is thus the only approach that delivers deterministic, and verifiable explanations.
@@ 6 Limitations
BELLA is a domain-independent method to produce deterministic post-hoc local explanations of tabular regression datasets. This scope entails some inherent limitations: First, as a generalist method, BELLA does not take into account domain-specific peculiarities, such as domain-specific weighting, dynamic scaling of feature distances, or domain-specific distance functions. Second, BELLA aims at simple and verifiable (and ultimately linear) explanations. With this, BELLA may fail to capture intricate, higher-order interactions of features. This problem is part of the intrinsic trade-off between simplicity and accuracy, which any post-hoc explainability approach finds itself in. Third, BELLA has to explore the whole neighborhood space without sampling, because it is deterministic. This incurs a high algorithmic cost, as our experiments show. This means that BELLA does not scale well to a large number of features, and the speed optimization of BELLA is left for future work.
There are also a number of ethical considerations: First, BELLA does not correct for biases in the data or the model. If the data contains biased samples, discriminating features, or an otherwise unprofessional selection of features or data points, these characteristics will be mirrored in BELLA’s explanations. We believe, however, that this is a feature rather than a bug: BELLA’s explanations make such biases visible to the user. It would be disastrous to correct them and thus convey the impression that the model does not have them. A second concern is the problem of overtrust, where users mistakenly take BELLA’s local explanation for a global correlation. That is an issue inherent to all local approaches, which would have to be countered by educating the user before using the approach. Finally, BELLA measures generality by the number of training examples that were considered to compute the explanation. While BELLA minimized the average error on this set of datapoints, it is possible that some examples of this neighborhood are not well explained.
@@ 7 Conclusion
We have presented BELLA, a deterministic approach to provide post-hoc local explanations for any regression black-box model, or indeed any static tabular dataset with a numeric variable to be explained. BELLA’s objective function ensures accurate, general, robust, and simple explanations. Detailed experiments show that BELLA outperforms state-of-the-art approaches on these desiderata, often by a wide margin.
Future work could investigate methods to improve the speed of BELLA. One possibility is to use other data structures to speed up the neighborhood search. Another possibility is to give up determinism and resort to sampling to further speed up the algorithm. Finally, one could investigate even if BELLA could replace black box models for making a prediction in the first place, following positive experiences with linear models elsewhere [Ismail et al., 2022]. We hope that our work can open the door to research along this line and others, and ultimately make machine learning models more interpretable.
@@ Appendix A Experiments with Random Forest as black-box model
Dataset LIME MAPLE BELLA SHAP
Auto MPG 1.57 $\pm$0.470 0.95 $\pm$0.272 1.52 $\pm$0.412 0.00
Bike 338.50 $\pm$20.92 61.23 $\pm$5.560 228.14 $\pm$14.48 0.00
Concrete 5.78 $\pm$0.844 2.22 $\pm$0.341 5.11 $\pm$0.681 0.00
Servo 0.45 $\pm$0.251 0.18 $\pm$0.152 0.45 $\pm$0.211 0.00
Electrical 0.01 $\pm$0.002 0.00 $\pm$0.001 0.01 $\pm$0.002 0.00
Supercond. 30.68 $\pm$1.407 0.96 $\pm$0.086 15.52 $\pm$0.481 0.00
White Wine 0.30 $\pm$0.022 0.17 $\pm$0.014 0.28 $\pm$0.023 0.00
Real Estate 5.19 $\pm$1.292 2.55 $\pm$0.962 5.04 $\pm$1.503 0.00
Wind 1.40 $\pm$0.197 0.70 $\pm$0.128 1.17 $\pm$0.184 0.00
CPU Activity 12.57 $\pm$1.110 0.63 $\pm$0.070 1.26 $\pm$0.010 0.00
Echocard. 3.31 $\pm$0.510 1.64 $\pm$0.246 3.21 $\pm$0.532 0.00
Iranian Churn 141.57 $\pm$20.44 9.46 $\pm$1.783 17.39 $\pm$2.923 0.00
Norm. avg. 0.07 $\pm$0.010 0.02 $\pm$0.004 0.05 $\pm$0.008 0.00
Table 8: Generality comparison (% - larger is better)